OpenStudy (anonymous):

Find the product. (y3)2 · y7

OpenStudy (anonymous):

we would use the rule that if (a^m)^n = a^mn and also that (a^m)*(a^n) = a^(m+n) so in this case we got (y^6)*y^7 and now we use the second rule because it has the same variable to obtain y^(6+7) = y^13 :)

OpenStudy (anonymous):

thats exactly what i thought it was, thanks :)

OpenStudy (anonymous):

awesomee